OpenStudy (anonymous):

Anders weighs 70.5 kg. He is carrying a bag with the mass of 7.5 kg och is going down in an elevator that is moving with a constant speed. a) How big is the force from the elevator floor to Anders? Suddenly the elevator starts to fall free b) how big is the force on Anders now? c) what happens to the bag if he drops it? How to approach b) and c)?

OpenStudy (shamim):

Total mass m=70.5+7.5=78kg Right?

OpenStudy (shamim):

Tell me weight w=?

OpenStudy (anonymous):

775N?

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

I wrote wrong in my question, he have the mass of 70 kg

OpenStudy (shamim):

If velocity is constant. Acceleration a=0 Right?

OpenStudy (shamim):

So totsl mass m=?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

U know weight w=mg

OpenStudy (anonymous):

But I know a), it is the same as his force but opposite direction! so its 775N

OpenStudy (shamim):

Both weight n force has same unit newton. Right?

OpenStudy (shamim):

So force n weight is equivalent

OpenStudy (anonymous):

775?

OpenStudy (shamim):

Andrew will apply a force on the floor of lift is its weight w

OpenStudy (anonymous):

Yes

OpenStudy (shamim):

Ya lift will apply equal n opposit force on andrew

OpenStudy (anonymous):

yes, thats what makes him dont go through the elevator floor

OpenStudy (shamim):

Make clearer plz. Mass m=?

OpenStudy (shamim):

m=70+7.5

OpenStudy (shamim):

Going to bathroom. Coming soon. Sorry

OpenStudy (anonymous):

70 kg + 7.5 kg = 77.5

OpenStudy (shamim):

So w=mg=77.5*9.8=?

OpenStudy (shamim):

Or u can use g=10m/s

OpenStudy (anonymous):

YES!!!! 775 N I have already said that? two times

OpenStudy (shamim):

R u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Shamim, I need help with question b and c, can you just write it in one message? not hundreds?

OpenStudy (anonymous):

@pompeii you there? need your help again :D

OpenStudy (shamim):

b) no force on andrew

OpenStudy (shamim):

c) the bag wil float

OpenStudy (anonymous):

Thanks shamim!

OpenStudy (shamim):

U r most welcome